# The unrealized potential of agroforestry for an emissions-intensive agricultural commodity

### v1

@@ Methods
Our study focused on the cocoa-growing regions of Ghana and Côte d’Ivoire, which together produce $\approx$60 % of the global cocoa supply [35] over an area of 7.04 million hectares [45]. In this region, closed tropical forests have largely been replaced by a mosaic of rotational fallows, slash-and-burn agriculture, and perennial cropping systems, with cocoa as one of the dominant crops. Cocoa farming in this region supports around two million farmers, most of whom manage small farms of three to five hectares [41, 22] and earn less than one dollar per day [57] but cocoa contributed to between 10 % and 20 % of the countries GDP. Despite its economic importance, the region faces declining productivity due to soil degradation, disease pressure, and threats from climate change, which are predicted to limit production in the future [73, 16, 36]. Furthermore, extensive deforestation has left little remaining primary forest [45, 67], making the biodiversity and carbon storage potential of cocoa agroforests increasingly vital.
@@
We used supervised machine learning to estimate shade-tree cover in cocoa production systems across Ghana and Côte d’Ivoire from optical satellite images, calibrated with ground-truth data. Below we describe the satellite data, ground-truth observations (sampling design and sampling methods), machine learning algorithms, and methods used to generate the final maps of shade-tree cover.
@@
Satellite data. We used satellite images from the Sentinel-2 constellation operated by the European Space Agency (https://sentinels.copernicus.eu/). Sentinel-2 satellites capture images of the earth across 13 spectral bands covering a wide range of the visible and near-infrared spectrum, at a spatial resolution between 10 and 60 meters per band, and at high revisit frequency. These images are ideal for our purpose as they are specifically collected to enable agro-environmental, forest, and vegetation mapping and monitoring. We collected atmospherically corrected, orthorectified surface reflectance images in 110 km × 110 km tiles across the study area. For each tile, we used the ten least cloud-affected images acquired between May 2020 and December 2022, and within 120 days of the the corresponding ground-truth data (described below).
@@
Sampling design for ground-truth data. To calibrate the estimation from satellite images (described above) we collected direct field observations of shade-tree cover. We achieved this by capturing imagery with drones that we flew directly above 827 individual cocoa farms, distributed within four agro-climatic zones covering Ghana and Côte d’Ivoire.
We focused our ground-truth sampling on farms that had been formally mapped in 2021 by two large cocoa trading companies. In Ghana, we sampled 698 farms distributed according to a four-level stratified sampling design: cocoa farms within communities within administrative districts within agro-climatic zones. More specifically, we first randomly selected eight administrative districts within each of four agro-climatic zones. The four agro-climatic zones were defined earlier by Bunn et al. [21], using a random forest machine learning algorithm to classify cocoa growing regions in Ghana according to climate and soil characteristics. Within each administrative district we randomly chose either one or two cocoa-growing communities, with the constraint that the selected communities have at least 25 recently mapped cocoa farms. We then randomly picked between 16 and 39 farms per community, ensuring that farms were distributed relatively evenly across the community, and that each individual farm was managed by a different farmer where possible.
In Côte d’Ivoire we collected ground-truth data at 60 additional farms. This was done in collaboration with a sampling program initiated by the Sustainable Cocoa Initiative Support Programme (SCISP) of the German Development Agency (GIZ). These farms were distributed in the Sud Comoé and Indénié-Djuablin administrative regions in the east and Cavally region in the west of Côte d’Ivoire. Subsequent to this, we extended the agro-climatic zoning methods of Bunn et al.  [21] to include Côte d’Ivoire (Supplementary methods, Fig. S1), which indicated that these 60 farms were distributed across two of the four original cocoa agro-climatic zones identified in Ghana, and an additional cool and wet zone in the west of Côte d’Ivoire that was not originally identified in Ghana, where it only occurs in the Volta region (Fig. S1). We calculate all results with respect to the updated zones across both countries. For effective ground-truthing of the satellite imagery, it was important that ground-truth data be collected on cocoa farms across the widest possible range of shade-tree cover levels that occur within the region. Because shade levels tended to be low in many of the farms we sampled according to our more formal sampling design (described above), we sampled an additional 69 farms in Ghana that we specifically targeted for their higher shade levels. To identify farms with higher cover before doing the on-ground measurements, we relied on farmer-reported shade-tree densities combined with interactive visual interpretation of Google Earth images. These additional high-cover farms were distributed in 30 cocoa-growing communities across four administrative districts in two agro-climatic zones.
@@
Sampling methods for ground-truth data. At each of the farms included in our ground-truth sampling design (described above), a DJI Mavic 2 drone was flown at a height of 80m to capture high-resolution images above the shade-tree canopy. All flights were planned, flown and processed using the cloud-based drone mapping software (DroneDeploy Inc., Version 2.236.0). For each farm we also had access to polygons delimiting its boundaries, so that we could ensure complete coverage of the farm area.
@@
For our final estimates of shade-tree cover in cocoa production systems, it was important to separate cover by shade trees from cover by cocoa trees. Therefore, we photogrammetrically reconstructed a digital surface model (DSM) of the canopy, and a digital terrain model (DTM), and subtracted the two to obtain a canopy height map for each farm. To determine a suitable height threshold to distinguish between shade trees and cocoa trees (which tend to be relatively short), we manually labelled 6800 polygons delimiting small sections of cocoa monocultures (total area 118 ha) in 230 cocoa farms distributed across the study area. These monoculture areas were identified using visual interpretation of orthophotos derived from the drone images. For all labelled polygons we then extracted the height of cocoa trees from the canopy height map. This analysis showed that 99.7% of all cocoa trees remained below a height of eight meters (Fig. S6), while previous work shows that heights of shade trees tend to be higher than eight meters [17]. We therefore used a threshold height of eight meters; vegetation above this height could be confidently attributed to shade trees rather than cocoa trees. This threshold height was used to isolate (‘mask’) cocoa canopies from shade-tree canopies in our canopy height maps computed from the drone data. The resulting shade cover maps served as local ground truth reference data to train, validate, and test the machine learning model for satellite data (described below), which was then deployed at country scale. All quality assessments of the resulting large-scale maps are also based on drone-derived reference data.
@@
Machine-learning methods and shade-cover model. We applied gradient boosting regression (GBR) [37] to estimate shade-tree cover from satellite images across Ghana and Côte d’Ivoire. GBR is an ensemble machine learning algorithm, renowned for its efficacy in predictive modeling tasks. It operates in a fully supervised setting, using the per-farm shade cover maps derived from drone images as target values and Sentinel-2 L2A data as input. GBR learns an ensemble model in sequential fashion, such that each ensemble member further decreases the overall prediction error via an additive correction. We employed the Python implementation of GBR in scikit-learn (https://scikit-learn.org/stable/), with the number of decision trees set to 2000 and the robust Huber loss [43] as objective function. For training, we enumerated all farm polygons that form the training portion of the dataset and computed the ground truth shade cover per $10\times 10$ meter Sentinel grid cell as the fraction of shade-tree pixels in the drone imagery that fall within that cell.
@@
In order to minimize the impact of temporal decorrelation caused by different acquisition dates of drone and satellite imagery, we selected the ten temporally closest satellite images within 120 days of each drone flight (after discarding overly cloudy images). We used all twelve channels of the Sentinel-2 L2A product, at the 10 meter grid spacing defined by the blue, green, red and near-infra red channels. Lower resolution channels were upsampled to the same 10 meter grid with nearest neighbor interpolation. Besides raw spectral intensities we also computed a set of vegetation-related indices from literature (NDVI, GRVI, RVI, GNDVI and NDMI) [1] and appended them as additional input channels, for a total of 17 input values per pixel. At each pixel we then concatenated the values of all neighbours within a centred 5$\times$5 neighbourhood (i.e., 50$\times$50 meters on the ground) to include contextual evidence and to avoid overly grainy results, for a total of 425 input features per pixel. These feature vectors were paired with the corresponding reference shade cover values, computed from the drone-based high-resolution maps.
@@
Feature vectors and associated reference values were extracted densely wherever ground truth values were available. One third of those values was set aside as independent test set, on the remaining two thirds we trained GBR regressors with five-fold cross validation. From the resulting five GBR models we kept only the best-performing one and evaluated it on the independent test set. All accuracy metrics in this manuscript were computed with that setting.
@@
Map of shade-tree cover. With the trained GBR model we estimated shade cover densely across the cocoa planted area of Ghana and Côte d’Ivoire (as per the cocoa map of Kalischek et al. [45]). To that end, we selected the ten least cloudy Sentinel-2 L2A images per tile between April and November 2022. All ten satellite images were processed into per-pixel feature vectors as derived above and mapped to per-pixel shade cover values with the GBR model, which yields ten shade cover maps. The ten estimates were averaged into a final shade cover map. Averaging across multiple acquisitions reduces the influence of remnant atmospheric effects on the final prediction. The dense shade cover map of Ghana and Côte d’Ivoire, with 10m GSD, is shown in Fig. 1.
@@
Data. AGBD estimates were derived from two key datasets. The first one were sparse, L4A footprint-level AGBD estimates (25 meter diameter) retrieved from spaceborne LiDAR (Light Detection and Ranging) with statistical models, as part of NASA’s Global Ecosystem Dynamics Investigation (GEDI) [29]. GEDI offers AGBD estimates with good geographic coverage, but only at sparse footprint locations (the dense L4B product, accumulated from L4A, has 1 km GSD, too coarse for our purposes [71]). By itself, the L4A product can therefore hardly be used to infer AGBD for the highly variable, small-scale cocoa farms distributed across our study area. Hence, we combined them with satellite-derived canopy height data (described further below). For our study, we acquired GEDI L4A AGBD measurements at individual 25 m diameter footprints distributed across the study area. Consistent with the timing of other data sources we used footprints acquired between April and November, 2022. We further filtered these to: 1) discard footprints that are deemed unreliable [50, 29, 28, 49], and 2) only use data from GEDI’s "power" beams, which are considered more reliable than the "coverage" beams for low biomass. After filtering, our AGBD dataset consisted of $\approx$2.1 million footprints with AGBD estimates, spread roughly evenly across the study area.
@@
The second key dataset we used are raster maps of vegetation height, with associated per-pixel uncertainty values, covering the entire study area. To obtain these maps we applied the global canopy height retrieval model of Lang et al. [50] to the ten least cloudy Sentinel-2 images per tile over the year 2022 and averaged the 10 retrieval results. Two properties make this dataset useful for our task of dense AGBD retrieval. First, canopy height is known to strongly correlate with AGBD  [51, 8], meaning that it is a suitable predictor where direct AGBD measurements are absent. Second, in contrast to the L4A product, the canopy height maps offer dense, wall-to-wall coverage, with only a few pixels rendered invalid due to persistent cloud cover. This extensive coverage can help interpolate between GEDI footprints, effectively ’filling in the gaps’ in that data.
@@
Deep-learning methods and biomass model. We used a convolutional neural network (CNN) to estimate AGBD across Ghana and Côte d’Ivoire. CNNs are a class of deep learning algorithms predominantly used for processing grid-like data, such as images. They apply a series of learnable filters to the input image, which are adept at capturing spatial patterns in a hierarchical fashion. For our work, we trained a CNN to densely predict AGBD at 50 meter grid spacing, using as input 150$\times$150 meter patches of canopy height (sampled at 10 meter grid spacing), and a periodic encoding of the path centre latitude. The latter makes it possible to also learn biomass patterns governed by the latitudinal location of the patch.
@@
[…]
Map of aboveground biomass. We generated a high-resolution (50$\times$50 m) map of AGBD across Ghana and Côte d’Ivoire, not limited to only the cocoa growing regions. To do so, we fed all 110 km × 110 km canopy height (and uncertainty) tiles of our study area through the trained CNN model. The resulting AGBD map is shown in Fig. 2.
Comparison of cocoa and forest areas. To compare area, carbon density, and total carbon storage across the three land-use classes — cocoa, disturbed forest, and undisturbed forest — we used the map of cocoa growing areas from Kalischek et al (2023)[45], while forest areas were derived from the Tropical Moist Forest (TMF) map produced by the European Commission’s Joint Research Centre (JRC), which depicts disturbed and undisturbed tropical moist forests for the year 2022 [84]. Both maps were bilinearly resampled to match the resolution of our AGBD map (50$\times$50 m), enabling us to filter it at its native resolution and calculate the area, mean and standard deviation of AGBD, as well as total aboveground biomass and 95% confidence intervals (CIs)for each land-use class. Note that areas mapped as cocoa by Kalischek et al. [45] were excluded from the forest map to ensure the land-use classes remained distinct. Biomass was converted to carbon assuming 47% of above-ground biomass is carbon, following Martin et al. (2011) [55].
@@
We found that shade-tree cover in cocoa production systems was relatively low, and well below several industry commitments (see Results section). This suggests that cocoa agroforestry systems may have substantial additional carbon sequestration potential. We estimated this potential under different future shade-tree cover scenarios.
Relationship between cover and biomass. To quantify how different levels of future potential shade-tree cover translates to carbon sequestration, one must first quantify the relationship between shade-tree cover and biomass. This was done by converting the AGBD estimates at each pixel into estimates of total aboveground biomass (AGB). To ensure consistent spatial alignment and enable accurate pixel-to-pixel comparisons, we bilinearly resampled the shade map (10$\times$10 meters) to match the resolution of the biomass map (50$\times$50 meters). We then quantified the relationship between shade-tree cover and total biomass with Bayesian linear regression, using our pixel-level estimates of shade-tree cover as the predictor variable, and our pixel-level estimates of total aboveground biomass as the response variable. Only pixels in cocoa growing areas with shade levels up to 40% cover were used for the analysis. That set is representative for the range of plausible agroforestry scenarios given current recommendations, and includes 99.89% of all data points. In addition to shade-tree cover, we included a second-order polynomial term (i.e., the predictor "shade-tree-cover2") to account for potential curvilinearity in the relationship between cover and biomass. We used a Gaussian likelihood function for the response variable, and for each model coefficient we used improper flat priors (a uniform distribution over the real numbers), a common uninformative prior in Bayesian analyses. The model was fitted using Stan [79] together with the R package brms[23]. To fit the model we ran four Markov chain Monte Carlo (MCMC) chains for 2000 iterations, with a warm-up period of 1000 iterations. This gave a total of 4000 samples. Trace plots were visually assessed to ensure model convergence, and we also checked that the scale reduction factor ($\hat{R}$) was approximately equal to one. We note that we did attempt to fit an error-in-variables model to account for measurement uncertainty in the predictor, but this model consistently failed to converge, so we proceeded with the simpler Bayesian regression model as described above. With respect to predicting carbon-sequestration potential the simpler model should be more conservative as, all else being equal, it will tend to underestimate the strength of the relationship between cover and biomass.
@@
Carbon stocks under different shade-tree cover scenarios. We used our empirical relationship between shade-tree cover and above-ground biomass stocks (Fig. S7) to estimate the amount of sequestered carbon under current conditions, and under four potential future levels of shade-tree cover (15%, 20%, 30% and 40%). These four threshold levels are indicative and were chosen because they correspond to Ghana’s national definition of forest [30], and to various industry recommendations for shade-tree cover in cocoa agroforests [72].

### v2

@@ Methods
Our study focused on the cocoa-producing landscapes of Ghana and Côte d’Ivoire, which span 7.04 million hectares [kalischek23natfood] and contribute roughly 60% of global cocoa production [FAO22]. Much of the region’s original closed tropical forest has been converted into a patchwork of land uses, including shifting cultivation, fallow systems, and perennial cropping systems—most notably cocoa. Cocoa farming in this region supports around two million farmers, most of whom manage small farms of three to five hectares [hainmueller2011sustainable, bymolt2018demystifying] and earn less than one dollar per day (Mighty Earth), even as cocoa contributes to between 10% and 20% of the countries’ GDP. Despite its economic importance, the region faces declining productivity due to soil degradation, disease pressure, and threats from climate change, which are predicted to limit production in the future [schroth2016b, blaser18, franzen2007]. Furthermore, extensive deforestation has left little remaining primary forest [kalischek23natfood, Renier_2023], underscoring the growing importance of cocoa agroforests for both biodiversity conservation and carbon sequestration.
@@
We used supervised machine learning to estimate shade-tree cover in cocoa production systems across Ghana and Côte d’Ivoire from optical satellite images, calibrated with ground-truth data. Below we describe the satellite data, ground-truth data, machine learning algorithms, and methods used to generate the final maps of shade-tree cover.
@@
Satellite data. We used satellite images from the Sentinel-2 constellation operated by the European Space Agency (https://sentinels.copernicus.eu/). Sentinel-2 satellites capture images of the earth across 13 spectral bands covering a wide range of the visible and near-infrared spectrum, at a spatial resolution between 10 and 60 meters per band, and at a revisit frequency of approximately 5 days for our study area. These images are ideal for our purpose as they are specifically collected to enable agro-environmental, forest, and vegetation mapping and monitoring. We collected atmospherically corrected, orthorectified surface reflectance images in $\times110$ km2 tiles across the study area.
For model building and evaluation, we aimed to collect a set of training images that were, (i) cloud-free, (ii) taken close to the time of ground-truth data collection, and (iii) representative of different imaging conditions. To do this we first downloaded Sentinel-2 images in four time intervals (March 2020–May 2021; August 2021-December 2021; April 2022-July 2022; August 2022-December 2022). We downloaded all Sentinel-2 revisits of a given tile that fell within these intervals. Within each interval we sorted the tiles by the image-level cloud-cover attribute and retained up to $N=20$ of the least cloudy acquisitions. Some farms lie within multiple tiles, thus increasing the total number of available images for these farms. The set of images was then passed on to the training phase, during which we discarded images with clouds over the ground-truth farms (based on the per-pixel cloud cover map) and images taken more than 120 days from the date of the ground-truth data collection. From the remaining images we retained the ten images taken temporally-closest to the date of the ground-truth data collection.
Our approach minimizes the impact of temporal decorrelation caused by different acquisition dates of ground-truth and satellite imagery, and also minimizes the effect of phenology on shade-cover estimates by using data (satellite and ground-truth) collected predominantly during the growing season when shade-tree cover is expected to be highest. Our learning approach should account for any remaining variation in shade-cover during this time, after which cover estimates (i.e., maps of shade-tree cover) are generated based only on satellite images taken during the 2022 growing season (see below). We excluded remaining cloud-affected pixels from the final selection of ten images using the Sentinel-2 pixel-level cloud mask. Because our learning approach should accommodate variation caused by cloud shadows, we did not mask for this attribute.
@@
Sampling design and methods for ground-truth data. To calibrate the estimation from satellite images we collected direct field observations of shade-tree cover. We achieved this by capturing imagery with drones that we flew directly above 827 individual cocoa farms, distributed across four agro-climatic zones covering Ghana and Côte d’Ivoire. We focused our ground-truth sampling on farms that had been mapped in 2021 by two large cocoa trading companies. We developed a new agro-climatic zoning scheme to stratify our ground-truth sampling across cocoa-growing regions in Ghana and Côte d’Ivoire (see Supplementary Methods). In Ghana, we sampled 698 farms distributed according to a four-level stratified sampling design: cocoa farms within communities within administrative districts within four agro-climatic zones (Extended Data Fig. Extended Data 1, Supplementary Methods). In Côte d’Ivoire we collected ground-truth data at 60 additional farms distributed across three agro-climatic zones (Extended Data Fig. Extended Data 1). For effective ground-truthing of the satellite imagery, it was important that ground-truth data be collected on cocoa farms across the widest possible range of shade-tree cover levels that occur within the region. Because shade levels tended to be low in general, we sampled an additional 69 farms that we specifically targeted for their higher shade levels (Supplementary Methods).
@@
For each farm, we flew a DJI Mavic 2 drone at a height of 80m to capture high-resolution images above the shade-tree canopy. All flights were planned, flown and processed using the cloud-based drone mapping software, DroneDeploy (Version 2.236.0). We used farm boundary polygons provided by cocoa trading companies to guide flight planning and ensure that each flight covered the area within or immediately surrounding the intended farm.
@@
For our final estimates of shade-tree cover it was important to separate cover from shade trees from cover from cocoa trees. We did this using tree height to distinguishing between shade trees and cocoa trees. To do this we first photogrammetrically reconstructed a digital surface model (DSM) of the canopy, and a digital terrain model (DTM), and subtracted the two to obtain a canopy height map for each farm (Supplementary Methods). We then used an empirically-validated threshold height of eight meters to distinguish between shade trees and cocoa trees (Supplementary Methods). This threshold height was used to isolate (‘mask’) cocoa canopies from shade-tree canopies in our canopy height maps computed from the drone data. The resulting shade-cover maps served as local ground-truth reference data to train, validate, and test the machine learning model for satellite data, which was then deployed at country scale. All quality assessments of the resulting large-scale maps are also based on drone-derived reference data.
@@
Machine-learning methods and shade-cover model. We applied gradient boosting regression (GBR) [friedman2001greedy] to estimate shade-tree cover from satellite images across Ghana and Côte d’Ivoire. GBR is an ensemble machine learning algorithm, renowned for its efficacy in predictive modeling tasks. It operates in a fully supervised setting, using the per-farm shade cover maps derived from drone images as target values and Sentinel-2 L2A data as input. GBR learns an ensemble model in sequential fashion, such that each ensemble member further decreases the overall prediction error via an additive correction. We employed the Python implementation of GBR in scikit-learn (version 1.2.1, https://scikit-learn.org/stable/), with the number of decision trees set to 2000 and the robust Huber loss [huber64] as objective function. For training, we enumerated all farm polygons that form the training portion of the dataset and computed the ground truth shade cover per $10\times 10$ meter Sentinel grid cell as the fraction of shade-tree pixels in the drone imagery that fall within that cell.
@@
We used all twelve channels of the Sentinel-2 L2A product, at the 10 meter grid spacing defined by the blue, green, red and near-infra red channels. Lower resolution channels were upsampled to the same 10 meter grid with nearest neighbor interpolation. Besides raw spectral intensities we also computed a set of vegetation-related indices from literature (NDVI, GRVI, RVI, GNDVI and NDMI) [abderazak96] and appended them as additional input channels, for a total of 17 input values per pixel. For each pixel we then concatenated the values of all neighbours within a centred $\times5$ neighbourhood (i.e., $\times50$ meters on the ground). This approach avoids overly grainy results, and, assuming that differences between a centre pixel and its neighborhood are not random, allows us to use features from surrounding pixels to better predict cover for centre pixels. We also note that this neighbourhood size was empirically validated using independent validation sets in order to identify the neighbourhood that provided the best predictions. Ultimately there were 425 input features per pixel. These feature vectors were paired with the corresponding reference shade cover values, computed from the drone-based high-resolution maps.
@@
Feature vectors and associated reference values were extracted for every pixel (i.e. densely) wherever ground-truth values were available. One third of those values was set aside as independent test set, on the remaining two thirds we trained GBR regressors with five-fold cross validation to identify optimal hyperparameters. A final model was trained on the entire dataset (excluding test set) and evaluated on the test set. All accuracy metrics in this manuscript were computed with that setting.
Map of shade-tree cover Using the trained GBR model, we estimated shade cover densely (i.e. per pixel) across the cocoa planted area of Ghana and Côte d’Ivoire, restricting predictions to areas mapped as cocoa-growing regions using the cocoa mask from Kalischek et al. (2023)[kalischek23natfood]. This map distinguishes shaded and unshaded cocoa farms, including young, mature, and old cocoa farms, from surrounding landscape features. We note that, despite some misclassification of minor (in terms of area) landscape elements such as roads and scattered farmhouses, the cocoa classification method of Kalischek et al. (2023)[kalischek23natfood] reliably identified cocoa farms regardless of their shade level, ensuring that both shaded and unshaded cocoa systems were included [Moraiti2024], which is a key requirement for meeting the goals of our study. To generate the shade-cover map, we used Sentinel-2 L2A images taken between April 2022 and November 2022, aligning with the growing season in this region. For each tile, images taken during this period were sorted quantitatively according to their cloud cover using the Sentinel-2 image-level cloud cover percentage attribute, after which the ten least cloud-affected images were selected. We excluded remaining cloud-affected pixels in these images from our analyses based on the Sentinel-2 pixel-level cloud mask.
@@
All ten satellite images were processed into per-pixel feature vectors as derived above and mapped to per-pixel shade cover values with the GBR model, yielding ten shade cover maps. The ten estimates were averaged into a final shade cover map. Averaging across multiple acquisitions reduces the influence of remnant atmospheric effects on the final prediction. The final shade cover map of Ghana and Côte d’Ivoire (Fig. 1), at 10 m ground sampling distance (GSD), was visualized in QGIS (v3.34.12[QGIS]).
@@
Data. AGBD estimates were derived from two key datasets. The first one were sparse, L4A footprint-level AGBD estimates (25 m diameter) retrieved from spaceborne LiDAR (Light Detection and Ranging) with statistical models, as part of NASA’s Global Ecosystem Dynamics Investigation (GEDI) [duncanson2022aboveground]. GEDI offers AGBD estimates with good geographic coverage, but only at sparse footprint locations (the dense L4B product, accumulated from L4A, has 1 km GSD, too coarse for our purposes [saarela2018generalized]). By itself, the L4A product can therefore hardly be used to infer AGBD for the highly variable, small-scale cocoa farms distributed across our study area. Hence, we combined them with satellite-derived canopy height data (described further below). For our study, we acquired GEDI L4A AGBD measurements at individual 25 m diameter footprints distributed across the study area. Consistent with the timing of other data sources we used footprints acquired between April and November, 2022. We further filtered these to: 1) discard footprints that are deemed unreliable [lang2023high, duncanson2022aboveground, dubayah2022gedi, lanfranchi22], and 2) only use data from GEDI’s "power" beams, which are considered more reliable than the "coverage" beams for low biomass. After filtering, our AGBD dataset consisted of $\approx$2.1 million footprints with AGBD estimates, spread roughly evenly across the study area.
@@
The second key dataset we used are raster maps of vegetation height, with associated per-pixel uncertainty values, covering the entire study area. To obtain these maps we applied the global canopy height retrieval model of Lang et al. (2023) [lang2023high] to the ten least cloudy Sentinel-2 images per tile over the year 2022 and averaged the 10 retrieval results. Two properties make this dataset useful for our task of dense AGBD retrieval. First, canopy height is known to strongly correlate with AGBD  [lefsky05, asner12], meaning that it is a suitable predictor where direct AGBD measurements are absent. Previous work has shown that predicting AGBD directly from a dense canopy height map gives satisfactory results[lanfranchi22]. Second, in contrast to the L4A product, the canopy height maps offer dense, wall-to-wall coverage, with only a few pixels rendered invalid due to persistent cloud cover. This extensive coverage can help interpolate between GEDI footprints, effectively ’filling in the gaps’ in that data. More generally, it is computationally more lightweight to directly lift the existing canopy height map to an AGBD map instead of estimating AGBD values from Sentinel-2 images, especially if local or regional maps are needed as in our case.
@@
Deep-learning methods and biomass model. We used a convolutional neural network (CNN) to estimate AGBD across Ghana and Côte d’Ivoire. CNNs are a type of deep learning model designed to interpret spatially organized data, such as image pixels arranged in a grid They apply a series of learnable filters to the input image, which are adept at capturing spatial patterns in a hierarchical fashion. For our work, we trained a CNN to densely predict AGBD at 50 m grid spacing, using as input $\times150$ m2 patches of canopy height (sampled at 10 m grid spacing) and a periodic encoding of the path centre latitude. The latter makes it possible to also learn biomass patterns governed by the latitudinal location of the patch. We note that we use larger grid spacing ($\times50$ m2) and context window ($\times150$ m2) for the AGBD model than our cover model because we use different analytical approaches (CNN vs. GBR) and different types of input data to model these different variables, and because empirical cross validation identified the context windows with highest predictive power in each case.
@@
Following standard practice we train our model in an iterative manner using stochastic gradient descent and back-propagation, where model parameters are updated at every step to increase the likelihood of the ground truth targets under the model’s output distribution. Each sample’s contribution to the overall error is reweighted according to the inverse square root of the frequency of the corresponding biomass value in the overall dataset, to reduce the bias towards the predominant AGBD values. We divide the dataset into 20 equidistant bins from 0 to 200 tonnes ha-1 and assign the same frequency to all samples in each bin. Our CNN has six layers, with kernel size $\times3$ and channel depths {16, 32, 64, 128, 128, 128}. We use the Adam optimizer [KingmaB14] with a base learning rate (step size) of $10^{-5}$. To estimate the predictive standard deviation, and to boost performance, we train an ensemble of five models with different random initialisations. Before training, the data is split into training, validation and testing portions, by dividing each Sentinel-2 tile into five longitudinal stripes of width 20 km and assigning three stripes to training, one to validation and one to testing. Optimisation was run until the validation performance no longer improved, and model performance was evaluated on the independent test set.
Map of aboveground biomass We generated a high-resolution ($\times50$ m2) map of AGBD across Ghana and Côte d’Ivoire. This map includes all cocoa growing areas and non-cocoa growing areas. To generate this map, we fed all 110 km × 110 km canopy height (and uncertainty) tiles of our study area through the trained CNN model. The resulting AGBD map Fig. 2, at 50 m ground GSD, was visualized in QGIS (v3.34.12[QGIS]).
Comparison of cocoa and forest areas. To compare area, carbon density, and total carbon storage across the three land-use classes — cocoa, disturbed forest, and undisturbed forest — we used our 2022 shade-cover map to delineate cocoa-growing areas. Forest areas were derived from the Tropical Moist Forest (TMF) map produced by the European Commission’s Joint Research Centre (JRC), which depicts disturbed and undisturbed tropical moist forests for the year 2022 [vancutsem2021]. Both maps were bilinearly resampled to match the resolution of our AGBD map ($\times50$ m2), enabling us to filter it at its native resolution and calculate the area, mean and standard deviation of AGBD, as well as total aboveground biomass and 95% confidence intervals (CIs) for each land-use class. Note that areas mapped as cocoa in our shade-cover map were excluded from the forest map to ensure the land-use classes remained distinct. Biomass was converted to carbon assuming 47% of above-ground biomass is carbon, following Martin et al. (2011) [martin2011].
@@
To quantify how different levels of future potential shade-tree cover translates to carbon sequestration, one must first quantify the relationship between shade-tree cover and biomass. This was done by converting the AGBD estimates at each pixel into estimates of total aboveground biomass (AGB) per pixel. To ensure consistent spatial alignment and enable accurate pixel-to-pixel comparisons, we bilinearly resampled the shade map ($\times10$ m2) to match the resolution of the biomass map ($\times50$ m2). We then quantified the relationship between shade-tree cover and total biomass with Bayesian linear regression, using our pixel-level estimates of shade-tree cover as the predictor variable, and our pixel-level estimates of total aboveground biomass as the response variable (Supplementary Methods).
@@
We used our empirical relationship between shade-tree cover and above-ground biomass stocks (Fig. S2) to estimate the amount of sequestered carbon under current conditions, and under four potential future levels of shade-tree cover (15%, 20%, 30% and 40%). These four threshold levels are indicative and were chosen because they correspond to Ghana’s national definition of forest, and to various industry recommendations for shade-tree cover in cocoa agroforests (Sanial et al. (2020)).